OpenStudy (anonymous):

What is the exact value of the expression square root of 18 plus square root of 48 minus square root of 50? Simplify completely.

OpenStudy (anonymous):

it's sqrt(48)*sqrt(50), and not sqrt(48)+sqrt(50), right?

OpenStudy (anonymous):

\[3\sqrt{2}+4\sqrt{3}+5\sqrt{2} = 4\sqrt{3}-2\sqrt{2}\]